40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 26 17:45:21 2014, Steve, Update, VAC, vac pressure rose to 1.3 mTorr noN2.pngbacktoVacNormal.png
    Reply  Sun Dec 28 07:29:16 2014, Steve, Update, VAC, vacuum is normal 3d.png
    Reply  Tue Feb 3 15:20:13 2015, Steve, Koji, Update, VAC, Pneumatic pressure is being read again N2pressureReadingIsBack.png
    Reply  Sun Mar 22 18:42:40 2015, Steve, Update, VAC, vac pressure rose to 1.2mTorr NoN2.pngPressureReadingWorks.png
       Reply  Mon Mar 23 10:36:55 2015, ericq, Update, VAC, Pressure watch script 
          Reply  Fri Apr 24 17:30:32 2015, Jenne, Update, VAC, Pressure watch script broken 
             Reply  Sat Apr 25 18:50:47 2015, ericq, Update, VAC, Pressure watch script broken 
                Reply  Sat Apr 25 22:17:49 2015, rana, Update, CDS, MXstream restart script working (beta) 
                   Reply  Tue May 19 17:03:34 2015, Koji, Update, CDS, MXstream restart script working (beta) 
                      Reply  Tue May 19 18:38:33 2015, rana, Update, CDS, MXstream restart script working (beta) 
          Reply  Mon Apr 27 14:37:55 2015, Steve, Update, VAC, N2 pneumatic pressure watch  
             Reply  Thu Apr 30 16:30:25 2015, Steve, Update, VAC, N2 pneumatic pressure watch set up 
          Reply  Tue Jan 31 14:14:29 2017, rana, Update, VAC, Pressure watch script 
       Reply  Sun May 15 19:48:08 2016, Steve, Update, VAC, run out of N2 noN2.png
          Reply  Tue Jun 14 10:13:34 2016, Steve, Update, VAC, temporary N2 supply line afterN2work.png
             Reply  Fri Jun 17 15:22:31 2016, gautam, Update, VAC, N2 supply line restored after retiling c0vac_06172016.png
Message ID: 11156     Entry time: Sun Mar 22 18:42:40 2015     In reply to: 10843     Reply to this: 11159   12117
Author: Steve 
Type: Update 
Category: VAC 
Subject: vac pressure rose to 1.2mTorr 

 

Quote:

We run out of N2 for the vacuum system. The pressure peaked at 1.3 mTorr with MC locked. V1 did not closed because the N2 pressure sensor failed.

We are back to vac normal. I will be here tomorrow to check on things.

We run out of N2 for the vacuum system 6 hrs ago. The pressure rose to 1.2 mTorr with V1 closed. The interlock worked! See Nirogen presure reading fixed at http://nodus.ligo.caltech.edu:8080/40m/10968

 

The vacuum interlock: Nitrogen pressure transducer is reading the pneumatic pressure continously at the pump spool and c1vac1 processing it. When it drops below 60 PSI it closes V1 gate valve and V4 & V5.  Gate valve V1 needs minimum 60 PSI to close. It is critical that V1 is closed before you run out of Nitrogen so the IFO pressure is contained.

 

IFO vacuum is back to Vac Normal. The MC is locked.

cc4 = 2E-6 Torr with VM1 open.

 

Daily N2 consumption measured to be 530PSI as 3 days on 3-27-2015 but note: it does vary !

I have seen it as high as 900 psi  The long term average ~750 psi

Attachment 1: NoN2.png  91 kB  Uploaded Sun Mar 22 19:43:54 2015  | Hide | Hide all
NoN2.png
Attachment 2: PressureReadingWorks.png  26 kB  Uploaded Mon Mar 23 11:41:09 2015  | Hide | Hide all
PressureReadingWorks.png
ELOG V3.1.3-